OpenStudy (anonymous):

The graph of which equation has a vertex of (3, 0)? y = (x + 3)^2 y = (x - 3)^2 + 3 y = 2(x + 3)^2 y = 2(x - 3)^2

OpenStudy (anonymous):

The answer is y = 2(x - 3)^2. Quadratic Form y=a(x-h)^2+k Vertex=(h,k)

OpenStudy (anonymous):

Given vertex is = ( 3 , 0) That is (h ,k) = (3,0) standard form of of parabola is y = ( x- h)^2 + k Plug (h , k) values in the parabola equation [ h=3,k=0] So we get y = (x - 3)^2 + 0 y = (x-3)^2